Meeting:2009/02/26

From SOWNWiki
Jump to: navigation, search

Meeting (to be) held on 2009/02/26 at 19:00 in Faculty Board Room

Previous meeting: 19 February 2009 19:00:00 Next meeting: 5 March 2009 19:00:00


Minutes

  • SOWN[at]HOME Nodes (daveruss)
    • Return of University road #250 (Davetaz)
      • No Progress
    • #269 and #271 resetupnoded using SjH's new Update Network Config tasklist and openvpn server configs updated to use udp, although #269 was apparently never changed.
    • The majority of the update config script works however snmp doesn't seem to work properly. Note that ipkg is now opkg and they've moved things around on their site.
    • #263 previously deployed at Shakespeare Ave North
      • No Progress on figuring out why its slowed down the host's network
    • New Deployments
      • Problems with Oxford Avenue Node (SjH) - Needs collecting (Addendum: node has now come up)
      • There has been a new deployment in Portswood, near the Esco Express (is up) and one in Archers road (not currently up but was only given out today).
    • Requests
      • No new requests, but there are still nodes to be collected.


  • PHP Code Cleanup
    • Problem logging in with new community accounts has been fixed. Look on the above page to find out what has been done as there have been some database changes. Solved by making the default value for userlevel 0 not null. Although removing user_level altogether and replacing with bannned should be the long term solution.
    • New login page (SjH) has been tidied, functionalised, documented and deployed. Seems to be working and user count is up. Authentication and login screen no logs errors to syslog.
    • Need Ajax on user selecting box
      • Implemented on mac_hiding page, node admin page and bowling website (different project just to prove our code is modular and portable). Done by daveruss and added to form_functions.inc.php
    • Finishing MAC hiding page and documenting (daveruss) has been done. See Documentation


  • Replacing cron jobs with ECS Webserver SVN
    • To be reviewed as there are very few cron jobs now.


  • Problem with mac address for an authenticated client not being pushed out to all nodes (daveruss)
    • Informal test carried out on Thursday (19/02/2009) allowed connection to a second node without reauthentication. Zepler node was checked and had picked up the MAC address also.


  • General Presentations about SOWN
    • Recovering presentations from Dell laptop with corrupt Windows partition (SjH)
    • No progress on either creating a general presentation or recovering old presentations


  • WiMAX
    • Campus Tour
      • Went ok. Links to photos have been sent but a page is needed which explains the photos. Morse to do that tomorrow morning.
    • Did bluenowhere send the final proposal to the vendor?
      • As yet unknown.


  • SOWN Core Network Update (Davetaz)
    • 2 servers have been requested. One of which will be the sown backup which needs to be installed by the 21st march.



  • Testing of IPv6 on SUSU
    • Is broken in SUSU, needs investigating.


  • 802.11r (Jos)
    • Appears to be working but for some things such as 802.11s however more tests need to be done to establish exactly what is going on.

More news next week.


  • SOWN VoIP (Inipairus)
    • Nothing major to report but Bekki has poking (not of the facebook variety)



  • SOWN-single Eduroam cloud (SjH and Davetaz and morse)
    • DEV has been set up to accept OpenVPN connections again, 200 and above.
    • (SjH) bricked his node and PC while investigating this so he is restoring them both so he can get back on track.


  • Link-Layer Filtering (formerly Captive-DNS -> munging) (morse)
    • (morse) has been playing, it works for forwarding but not input or output. The solution is to add these hooks to the driver.


  • AOB
    • Ajax IRC logs have been done for the benefit of Dave's other half.
Facts about "2009/02/26"
Has date26 February 2009 19:00:00 +
Has end date26 February 2009 20:00:00 +
Has location32/4073 +